Find the size of the force that it would require to bring a 4000 kg roller coaster car from a velocity of +30.0 m/s to a velocity of -15.0 m/s in a time of 6.0 seconds.


1
Expert's answer
2021-03-23T11:10:21-0400

Let's first find the acceleration of the roller coaster:


"a=\dfrac{v_f-v_i}{t}=\dfrac{-15\ \dfrac{m}{s}-30\ \dfrac{m}{s}}{6\ s}=-7.5\ \dfrac{m}{s^2}."

Finally, we can find the force:


"F=ma=4000\ kg\cdot(-7.5\ \dfrac{m}{s^2})=-3.0\cdot10^4\ N."